### v3

# Einstein Telescope: Detection of Binary Black Hole Gravitational Wave Signals Using Deep Learning

Abstract
Expanding upon our prior work (Alhassan et al., 2022), where we evaluated the performance of single sub-detector data (SSDD) from the Einstein Telescope (ET) for binary black hole (BBH) detection using deep learning (DL). In this study, we explore the detection efficiency of BBHs using data combined from all three proposed sub-detectors of ET (TSDCD), employing five different lower frequency cutoffs ($F_{\text{low}}$): 5 Hz, 10 Hz, 15 Hz, 20 Hz, and 30 Hz, while maintaining the same match-filter Signal-to-Noise Ratio ($MSNR$) ranges as in our previous work: 4-5, 5-6, 6-7, 7-8, and >8. The Deep Residual Neural Network model (ResNet) was trained and evaluated for the detection of BBH gravitational wave signals using both TSDCD and SSDD. Compared to SSDD, the detection accuracy from TSDCD has shown substantial improvements, increasing from $60\%$, $60.5\%$, $84.5\%$, $94.5\%$ to $78.5\%$, $84\%$, $99.5\%$, $100\%$, and $100\%$ for sources with $MSNR$ of 4-5, 5-6, 6-7, 7-8, and >8, respectively. In a qualitative evaluation, the ResNet model detected sources at 86.601 Gpc, with an averaged $MSNR$ of 3.9 (averaged across the three sub-detectors) and a chirp mass of 13.632 at 5 Hz. The results demonstrate a notable accuracy improvement for lower $MSNR$ ranges (4-5, 5-6, 6-7) by $18.5\%$, $24.5\%$, and $13\%$, respectively, and by $5.5\%$ and $1.5\%$ for higher $MSNR$ ranges (7-8 and >8). TSDCD proves suitable for near-real-time detection and can benefit from a more powerful setup.
Gravitational waves(678) — Gravitational wave detectors(676) — Astronomy data analysis(1858)
software: LALSuite (LIGO Scientific Collaboration et al., 2018), pyCBC (Nitz et al., 2023)
1 Introduction
This is a continuation of our previous work (Alhassan et al., 2022) (WTM1 hereafter referred to as) on the detection of binary black holes (BBHs) gravitational wave (GWs) signals from the Einstein Telescope (ET) using deep learning (DL). ET (Punturo et al., 2010; Abernathy et al., 2011; Maggiore et al., 2020) is designed [^1] See https://www.et-gw.eu/index.php/relevant-et-documents for relevant documents on the ET design. to be an underground GWs detector, where the seismic noise is much lower, and hence the level of Newtonian noise. ET as shown in Figure 1, will consist of three nested arms (hereafter referred to as sub-detectors) with 10 Km long each, in an equilateral triangle shape, with arm-opening angles of $60^{\circ}$ (Hild et al., 2009; Branchesi et al., 2023). Compared to 3 km length for advanced Virgo (Acernese et al., 2014) and KAGRA (Aso et al., 2013; Akutsu et al., 2020), and 4 km for advanced LIGO (Collaboration et al., 2015), the ET 10 km V-shaped sub-detectors length will significantly improve its sensitivity allowing the observation of GWs at lower frequency that reach $\approx 2$ Hz (Hild et al., 2011). Each sub-detector of ET will consist of two interferometers, one optimised for low-frequency and one for high-frequency sensitivity. In this work, it is assumed that each sub-detector has a single interferometer.
Figure 1: Simplified layout of the three core sub-detectors of the ET telescope.
The current standard method for detecting GW signals is match-filtering, considered the most sensitive algorithm targeting compact binary sources, such as BBHs with spin-aligned components on quasi-circular orbits (Gerosa et al., 2013; Rodriguez et al., 2016, 2015). One of the main drawbacks of such algorithms is their computational complexity, especially when targeting eccentric BBHs systems (Harry et al., 2016). In addition, match-filtering algorithms might miss signals generated by compact binary populations in dense stellar environments (Huerta et al., 2014; Klimenko et al., 2016; Huerta et al., 2017). These limitations open the door to potential applications of deep learning algorithms, which do not require waveform templates for the detection.
The application examples of deep learning in astrophysics, and more specifically in GWs, were shown in WTM1. Notably, more work has been presented since then, highlighting the significant interest in this promising technology for addressing the considerable challenge posed by the vast amount of data expected from current and upcoming detectors. The Spy project (Glanzer et al., 2023) applied machine learning for glitch classification using data up to the end of the third observing run of LIGO. In this analysis, 233,981 glitches from LIGO Hanford and 379,805 glitches from LIGO Livingston were classified into morphological classes. A related study employed generative adversarial networks (GANs) to simulate hundreds of synthetic images representing the 22 most common types of glitches observed in the LIGO, KAGRA, and Virgo detectors. A neural network model was then used to classify the simulated glitches, achieving an average accuracy of $99.0\%$ (Powell et al., 2023). Accelerating Artificial Intelligence models for rapid GW detection was recently conducted by Chaturvedi et al. (2022). In thier work, the ThetaGPU supercomputer was leveraged with the NVIDIA TensorRT-optimized AI ensemble to process a month of advanced LIGO data; the process took 50 seconds. In Yan et al. (2022), a novel model called a gravitational wave squeeze-and-excitation shrinkage network (GW-SESNet) was inspired by attention mechanisms, a mechanism in deep learning used to highlight important features, for the classification of real signals, noise, and glitches. GW-SESNet utilized the coherent signal-to-noise ratio (SNR) from multiple detectors and achieved recall values of 0.84, 0.98, and 0.98 for noise, real signals, and glitches, respectively. The role of advanced machine learning tools in the coming decades, either independently or in tandem with traditional data analysis techniques, may become crucial considering the vast amount of the data expected (Shah et al., 2023).
The detection of GW sources represents the first step in gravitational wave (GW) data analysis; hence, quick and fast detection is required for subsequent steps, such as source parameter estimation or alerting other telescopes for multi-messenger follow-up and studies of the detected source. Fast detection can also be employed for data marking, facilitating a detailed Bayesian parameter estimation analysis of the relevant part of the detector output at a later stage. Coherent multi-detector observation, in which multiple detectors are coupled together, is beneficial for improving the coherent SNR (Macleod et al., 2016), making it a current trend in gravitational wave signal detection.
In WTM1, our work focused on the detection of BBHs sources that were generated using only one single sub-detector of ET, with low-frequency-cutoff (starting frequency of the binary systems, which increases as the binary system evolves and the components get closer to each other) of 30 Hz, and with five match-filter SNR (hereafter referred to as $MSNR$) ranges: 4-5, 5-6, 6-7, 7-8 and >8. Expanding upon our previous work, we explore the detection efficiency of BBHs using data combined form all the three proposed sub-detectors of ET (hereafter referred to as three sub-detectors coherent data (TSDCD)), with five different lower frequency cutoff ($F_{low}$): 5 Hz, 10 Hz, 15 Hz, 20 Hz and 30 Hz, employing identical $MSNR$ ranges for each frequency. In addition, we used the same luminosity distance ($D_{L}$) range of 145 Mpc - 120 Gpc as in WTM1 for our BBH sources. The main objective is to compare the previous obtained results from a single sub-detector data (hereafter referred to as SSDD), to the results from TSDCD.
The structure of this work goes as follow: in Section 2 we describe the method and tools were used for the generation of TSDCD and the prepossessing pipeline employed for spectrogams preparation. Section 3 discusses Deep Residual Neural Networks, and the training and evaluation processes. Evaluating ResNet model on the TSDCD is presented in section 4 and 5. Finally, a summary of this work and a future work plans is presented in Section 6.
2 ET data simulation
The strain in each sub-detector of ET $h(t)_{E\alpha}$ is given as :
$h(t)_{E\alpha}=\frac{1}{4}F_{\alpha}^{+}(\theta,\phi,\psi)h_{+}(t)+F_{\alpha}^{\times}(\theta,\phi,\psi)h_{\times}(t),$ (1)
where $\alpha$ is the index of the sub-detector; $\alpha=(1,2,3)$. $F_{\alpha}^{+}$ and $F_{\alpha}^{\times}$ are antenna response functions for the two GW polarizations of each sub-detector $\alpha$, which depend on the sky localization (right ascension $\theta$ and declination $\phi$), polarization angle $\psi$, and take into account the motion of the Earth (Jaranowski et al., 1998). For a single sub-detector, $F^{+}$ and $F^{\times}$ at time $t$ and an angel $\eta$ between the interferometer arms are given as:
$F_{+}(t)$ $=\sin\eta[a(t)\cos 2\psi+b(t)\sin 2\psi],$ (2)
$F_{\times}(t)$ $=\sin\eta[b(t)\cos 2\psi-a(t)\sin 2\psi],$ (3)
where:
$a(t)$ $=\frac{1}{16}\sin 2\gamma(3-\cos 2\lambda)(3-\cos 2\delta)\cos[2(\alpha-\phi_{r}-\Omega_{r}t)]$
$\quad-\frac{1}{4}\cos 2\gamma\sin\lambda(3-\cos 2\delta)\sin[2(\alpha-\phi_{r}-\Omega_{r}t)]$
$\quad+\frac{1}{4}\sin 2\gamma\sin 2\lambda\sin 2\delta\cos[\alpha-\phi_{r}-\Omega_{r}t]$
$\quad-\frac{1}{2}\cos 2\gamma\cos\lambda\sin 2\delta\sin[\alpha-\phi_{r}-\Omega_{r}t]$
$\quad+\frac{3}{4}\sin 2\gamma\cos 2\lambda\cos 2\delta,$ (4)
$b(t)$ $=\cos 2\gamma\sin\lambda\sin\delta\cos[2(\alpha-\phi_{r}-\Omega_{r}t)]$
$\quad+\frac{1}{4}\sin 2\gamma(3-\cos 2\lambda)\sin\delta\sin[2(\alpha-\phi_{r}-\Omega_{r}t)]$
$\quad+\cos 2\gamma\cos\lambda\cos\delta\cos[\alpha-\phi_{r}-\Omega_{r}t]$
$\quad+\frac{1}{2}\sin 2\gamma\sin 2\lambda\cos\delta\sin[\alpha-\phi_{r}-\Omega_{r}t],$ (5)
where $\lambda$ denotes the latitude of ET’s location, which we defined same as VIRGO. $\Omega_{r}$ is Earth’s rotational angular velocity, and $\phi_{r}$ is the phase defining the position of the Earth in its diurnal motion at $t=0$.
$\gamma$ determines the orientation of the detector arms and is measured counter-clockwise from the east to the bisector of the interferometer arms. For the ET in the triangular configuration, $\eta=60^{\circ}$.
The polarizations of the GW $h_{+}(t)$ and $h_{\times}(t)$ in Eq (1), from an inspiraling compact binary system such as BBH, are given as:
$h_{+}(t)$ $=-\frac{1+\cos^{2}\iota}{2}\left(\frac{GMc^{2}D_{L}}{2(tc-t)}\right)^{1/4}\times\cos\left[2\Phi_{c}+2\Phi(t-tc;M,\mu)\right],$ (6)
$h_{\times}(t)$ $=-\cos\iota\left(\frac{GMc^{2}D_{L}}{2(tc-t)}\right)^{1/4}\times\sin\left[2\Phi_{c}+2\Phi(t-tc;M,\mu)\right],$ (7)
where $\iota$ denotes inclination angel of the orbital plane of the binary system relative to the observer. $c$ and $G$ represent the speed of light and the gravitational constant respectively. The quantity $\mu$ denotes the binary system’s reduced mass. $t_{c}$ and $\Phi_{c}$ represent, respectively, the time and the phase at which the waveform terminates. $t-t_{c}$ represents the time difference between the current time $t$ and the coalescence time $t_{c}$. The function $\Phi(t-t_{c};M,\mu)$ represents the binary system’s orbital phase. $M$ is the source chirp mass, which depend on the mass component of the binary system $m_{1}$ and $m_{2}$, and is given as:
$M=(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$ (8)
In WTM1, our emphasis was on simulating BBH sources using a single sub-detector of ET. In this work, we replicated the exact BBH source parameters, data generation tools, and methods employed in WTM1 (for details, please see Alhassan et al., 2022, section 2) for each sub-detector of ET to obtain TSDCD.
2.1 Simulation
We obtained the mass components and redshift ($z$) for simulating BBH systems from (Belczynski et al., 2020). A broad set of compact binary models was developed by Belczynski et al. (2020) using the population synthesis code StarTrack (Belczynski et al., 2002b, c, a), which has been significantly upgraded in Belczynski et al. (2008) to focus on processes leading to the formation and further evolution of compact objects.
Restricting the BBH masses to 15-56 $M_{\odot}$ and redshift $z$ values from 0.033 to 11 allows us to focus on low and medium masses within our selected MSNR ranges. In addition, we have considered five $F_{\text{low}}$ bounds, namely 5 Hz, 10 Hz, 15 Hz, 20 Hz, and 30 Hz. Other parameters such as: the sky position angles ($\theta$ and $\phi$), $\iota$, $\psi$, and coalescence phases, were sampled from a uniform distribution as in WTM1.
Assuming the standard $\Lambda$CDM cosmology (Planck Collaboration et al., 2016), where the Hubble constant $H_{0}=67.74\,\text{km s}^{-1}\,\text{Mpc}^{-1}$, energy density parameter $\Omega_{\Lambda}=0.6910$, and matter density parameter $\Omega_{m}=0.3075$. We convert the redshift $z$ to $D_{L}$ using the analytical approximation provided by Adachi & Kasai (2012) for the $D_{L}$ between the BBH source and the detector, given as:
$D_{L}=\left(1+z\right)cH_{0}\int_{0}^{z}\frac{dz}{\sqrt{\Omega_{\Lambda}+\Omega_{m}(1+z)^{3}}}$ (9)
where $c$ is the speed of light in the vacuum.
We utilized the latest time-domain phenomenological higher-order mode model IMRPhenomD (Pürrer et al., 2023), which is part of LALSuite (LIGO Scientific Collaboration et al., 2018), employing its Python interface (Wette, 2020). IMRPhenomD is designed for generating GW signals from merging BBH with non-precessing spins. For each sub-detector, using the parameters above, we individually generated a BBH GW signal ($h(t)_{E1},h(t)_{E2},h(t)_{E3}$), then added them into the simulated Gaussian detector noise $n$, assuming an additive, Gaussian and stationary noise model as illustrated in Figure 2. This process results in two datasets: The data $d$ (if the signal $h$ is present) and $d_{n}$ (if the data contains noise only), can be written as follow:
$d(t)$ $=h(t)_{E1}+h(t)_{E2}+h(t)_{E3}+n(t)$ (10)
$d_{n}(t)$ $=n(t)$ (11)
Figure 2: A 5-second segment of simulated data for all ET sub-detectors. The gray line represents detector noise $n(t)$. The blue, green, and red lines represent the injected BBH signals from E1, E2, and E3, respectively. The injected BBH source has an MSNR in the range $7-8$ with component masses $m_{1}=20\,M_{\odot}$ and $m_{2}=25\,M_{\odot}$, and $F_{\text{low}}$ of 30 Hz.
The improvement in SNR through the utilization of data from all sub-detectors of ET combined can be assessed by calculating the coherent SNR. Coherent SNR $\text{SNR}_{\text{coh}}$ enhances the observed signal’s SNR by incorporating the coherent combination of multiple detectors (Macleod et al., 2016). To quantify the improvement in MSNR resulting from the combination of each sub-detector, we calculated $\text{SNR}_{\text{coh}}$ using the following formula:
$\text{SNR}_{\text{coh}}=\sqrt{\sum_{\alpha=1}^{3}\text{MSNR}_{\alpha}^{2}}$ (12)
Figure 3: $\text{SNR}_{\text{coh}}$ from the combination of each sub-detector.
2.2 Spectrograms
The Short Time Fourier Transform (STFT) (Yin et al., 2013) was used to generate our TSDCD spectrogram samples ($S_{E123}$), as shown in Figure 4. Three spectrograms ($S_{E1}$, $S_{E2}$, and $S_{E3}$) were produced for each source in each sub-detector ($E1$, $E2$, and $E3$), and then stacked to form an RGB image $S_{E123}$, represented as:
$S_{E123}=\begin{bmatrix}R(i,j)\\ G(i,j)\\ B(i,j)\\ \end{bmatrix}$ (13)
Where $R(i,j)$, $G(i,j)$ and $B(i,j)$ are the normalized intensity values for each channel at pixel $i,j$ in $S_{E1}(i,j)$, $S_{E2}(i,j)$, and $S_{E3}(i,j)$, calculated as:
$R(i,j)$ $=\frac{S_{E1}(i,j)}{\text{max\_intensity}}$ (14)
$G(i,j)$ $=\frac{S_{E2}(i,j)}{\text{max\_intensity}}$ (15)
$B(i,j)$ $=\frac{S_{E3}(i,j)}{\text{max\_intensity}}$ (16)
where max_intensity represents the maximum intensity value across all $S_{E1}$, $S_{E2}$, and $S_{E3}$. A total of 125,000 sources were generated, with each of the five defined $F_{low}$ groups containing 25,000 sources. These sources were evenly distributed among our five $MSNR$ intervals. It is important to note that $MSNR$ values vary between the three detectors using the same parameters due to orientation and polarization. For each source, at least one detector must have the correct $MSNR$ value to be accepted into the $MSNR$ interval.
Similarly, to generate only-noise spectrograms, the following steps were taken: 1) A Gaussian noise with a color matching the power spectrum density of ET, sampled at ET’s frequency, was generated. 2) Three random segments of 5 seconds each were selected. 3) STFT was performed on each segment to produce a spectrogram. 4) All three only-noise spectrograms were stacked together.
Figure 4: Generation and processing of the spectrograms from TSDCD
3 ResNet: Deep Residual Neural Networks
In WTM1, the performance of the most popular five computer vision neural networks models were evaluated for the detection of BBH GW signals, namely VGG-16 and VGG-19 (Simonyan & Zisserman, 2014), DenseNet (Huang et al., 2016), ResNet (He et al., 2015; Wu et al., 2018). The ResNet model showed the best overall performance, and thus, our goal in this study is to assess its performance on TSDCD compared to SSDD.
One of the great advantages of ResNet network, is its ability to overcome the vanishing gradient problem, which considered as one of the challenging issues in training deep and complex neural networks models (Hu et al., 2018; Roodschild et al., 2020). This was achieved by introducing residual blocks, which enable the training of hundreds of layers (Goceri, 2019a, b; Hasan et al., 2021). These blocks use an additive attribute to merge previous and future layers. ResNet is composed of several residual blocks. Each residual block is formed by adding a residual connection every two layers of conventional convolution. The key concept behind the residual connection, which is considered a pivotal breakthrough in deep learning, is to learn a residual function capable of representing the difference between the input $X$ and the target $\mathcal{F}(X)$ (the desired mapping learned by the layer). A residual mapping $\mathcal{R}(X)$ is learned as follows:
$\mathcal{R}(X)=\mathcal{F}(X)-X.$ (17)
Then the sum of the residual mapping and the input is the output of the layer ($Y$) and can be calculated as:
$Y=X+\mathcal{R}(X).$ (18)
Due to this connection, the original output will be superimposed before sending it to the future layer. And hence, mitigate and prevent any gradient explosion or disappearance when using the backpropagation (Munro, 2010, for details about backpropagation please see), and allow for training of very deep networks. As previously mentioned in WTM1, ResNet is available in various architectures, distinguished by the number of layers. This includes ResNet-50 and ResNet-101, which comprise 50 and 101 convolutional layers, respectively. We adopt the ResNet-101 architecture (hereafter referred to as ResNet) in this work, as was previously done in WTM1.
3.1 ResNet training
We kept the same settings for training the ResNet model as in our previous work (WTM1). However, the dataset size has increased fivefold in TSDCD compared to the single sub-detector data (SSDD) we used before. This expansion is a result of incorporating five different $F_{low}$ values. In machine learning (ML), traditionally, the data is split into three different sets for training, testing, and evaluating the model (Murphy, 2012; Raykar & Saha, 2015; Medar et al., 2017; Nguyen et al., 2021; Tan et al., 2021).
Table 1 displays the total number of TSDCD samples utilized for ResNet model training, testing, and validation. All $F_{low}$ values and $MSNR$ ranges are evenly represented in the randomly chosen testing and validation sources, to ensure an unbiased evaluation of the model’s performance. The three sub-detector combined spectrograms consist of three layers, where each layer belongs to one sud-detector. Hence, the input layer of ResNet was adjusted to a shape size of 365 × 42 × 3, which describes the width, height, and number of layers of the input image, respectively. A batch size of 256, learning rate of 0.0001, and RMSprop optimizer (Vijendra Babu et al., 2020; Xu et al., 2021; Elshamy et al., 2023) were used for training the ResNet model for 200 epochs.
Type Number of Sample Train Test Val
Injected 125,000 85,000 20,000 20,000
Only noise 125,000 85,000 20,000 20,000
Total 250,000 170,000 40,000 40,000
Table 1: Total number of TSDCD spectrograms: injected and only-noise spectrograms for training, testing and validation.
4 TSDCD evaluation
Figure 5: $MSNR$ versus detection’s accuracy, Top: SSDD (from WTM1) versus TSDCD for sources with $F_{low}$ of 30 Hz. Bottom: TSDCD for sources with $F_{low}$ of 5 Hz, 10 Hz, 15 Hz, 20 Hz and 30 Hz.
Figure 6: Confusion matrix comparison for $MSNR$ in the ranges 4-7 between SSDD and TSDCD, with $F_{low}$ of 30 Hz.
Figure 7: Confusion matrix comparison for $MSNR$ 7-8 and greater between SSDD and TSDCD, with $F_{low}$ of 30 Hz.
To evaluate the performance of ResNet on the TSDCD, similar to WTM1, we generated a classification report displaying precision, recall, and F1-score values (Bishop, 2006) for our testing dataset, as shown in Table 2. Recall, also known as sensitivity (Wang et al., 2020), represents the quantity of correctly classified and misclassified sources, where ResNet achieved $80\%$ and $99\%$ on injected and only-noise spectrograms. This result demonstrates a notable trend, indicating that only-noise spectrograms are highly unlikely to be misclassified as injected ones across all the different $F_{low}$ and $MSNR$ ranges. Precision, assessing the quality of classification, achieved an average of $90\%$ by ResNet. The overall performance of ResNet, measured by the F1-score — a harmonic mean of precision and recall — also attained an average of $90\%$.
Type precision recall f1-score support
Injected 0.993 0.818 0.897 20,000
Only noise 0.845 0.994 0.914 20,000
avg/ total 0.919 0.906 0.905 40,000
Table 2: Classification report showing the precision, recall and F1-score metrics on the testing dataset.
To compare the TSDCD results with those obtained in our prior investigations in WTM1, we generated a TSDCD with an $F_{low}$ of 30 Hz only with all 5 $MSNR$ ranges. This dataset consists of 4000 sources, created using identical parameters to the testing dataset in WTM1. In Figure 5, the top panel displays the accuracy of the detection acquired for the TSDCD, alongside the accuracy obtained by SSDD in WTM1. The plot clearly demonstrates a significant improvement across all five $MSNR$ ranges, particularly at lower $MSNR$ values. The detection accuracy has improved from $60\%$, $60.5\%$, $84.5\%$, $94.5\%$, and $98.5\%$ to $78.5\%$, $84\%$, $99.5\%$, $100\%$, and $100\%$ for sources with $MSNR$ of 4-5, 5-6, 6-7, 7-8, and greater than 8, respectively. This shows a significant improvement of approximately $30.83\%$, $39.00\%$, $17.75\%$, $5.81\%$ and $1.52\%$ for sources with $MSNR$ of 4-5, 5-6, 6-7, 7-8, and greater than 8, respectively. It is essential to note that achieving $100\%$ accuracy for sources with $MSNR$ greater than 8 means the successful detection of all sources within this specific range.
The bottom panel in Figure 5 displays the detection accuracy from our testing dataset, considering each $F_{low}$ individually. The results indicate a generally similar performance across all frequencies and $MSNR$ ranges, with relatively inferior performance observed for lower $MSNR$ values (4-6) at $F_{low}$ of 5 Hz, as expected.
In a quantitative evaluation, we compared the number of misclassified injected and only-noise spectrograms using confusion matrices (Stehman, 1997). We used 200 randomly selected samples (100 samples from each class) from each $MSNR$ range, with an $F_{low}$ of 30 Hz, obtained from both SSDD and TSDCD to generate the confusion matrices. As shown if Figure 6, a significant decrease can be observed in the number of misclassified Injected spectrograms from 66, 54, 8, and 1 to 50, 25, 0, and 0 for $MSNR$ 4-5, 5-6, 6-7, and 7-8, respectively, on the TSDCD, in contrast to the results from SSDD. For only-noise spectrograms, in TSDCD, only tow samples were misclassified as injected, in contrast to the 19 total misclassifications observed in SSDD. In Figure 7, the confusion matrices for sources with $MSNR$ greater than 8 illustrate cases where no only-noise spectrogram was identified as injected in TSDCD, in contrast to 3 in SSDD.
5 Evaluation on ET’s mock data
As previously done in WTM1, three sub-detectors combined mock data was generated to assess the efficiency of utilizing the TSDCD for near-real-time detection and to compare it with previously obtained results from SSDD. Five hours and 50 seconds of time-series data, encompassing all $MSNR$ ranges, were generated for each $F_{low}$ (5 Hz, 10 Hz, 15 Hz, 20 Hz, and 30 Hz) for each detector. At every five-second interval in this time-series data, a signal is injected at a random position, resulting in a total of 3,610 injected sources, with each $MSNR$ range containing 722 sources. The total duration of the data for each detector is 25 hours, 4 minutes, and 16 seconds. To compare with our previously obtained results, we utilize the data with an $F_{low}$ of 30 Hz.
Figure 8: Inferencing on ET’s mock data
To perform detection on this time-series data, we follow four main steps, as illustrated in Figure 8: 1) slide a 5-second window simultaneously across the datasets of the three detectors, 2) generate a spectrogram for each window, 3) crop and stack spectrograms, and 4) feed the stacked spectrogram into the trained ResNet model for prediction.
The False Positive Rate (FPR) (Burke et al., 1988) quantifies how often the model incorrectly classifies only-noise samples as injected ones, defined as follows:
$FPR=\frac{FP}{FP+TN}$ (19)
Where FP is the number of false positives, and TN is the number of true negatives.
Figure 9: Top: A comparison of FAR values between SSDD (from WTM1) and TSDCD for sources with $F_{low}$ set at 30 Hz. Bottom: FAR values exclusively from TSDCD for sources with $F_{low}$ at 5 Hz, 10 Hz, 15 Hz, 20 Hz, and 30 Hz.
Figure 10: Right: $MSNR$ of all undetected BBH sources from SSDD (from WTM1) and TSDCD. Left: $MSNR$ of undetected sources exclusively from TSDCD for sources with $F_{low}$ at 5 Hz, 10 Hz, 15 Hz, 20 Hz, and 30 Hz.
Figure 11: Right: Chirp mass $M$ of all undetected BBH sources from SSDD (from WTM1) and TSDCD. Left: Chirp mass $M$ of undetected sources exclusively from TSDCD for sources with $F_{low}$ at 5 Hz, 10 Hz, 15 Hz, 20 Hz, and 30 Hz.
Figure 12: Right: $D_{L}$ of all undetected BBH sources from SSDD (from WTM1) and TSDCD limited to 70 Gpc. Left: $D_{L}$ of undetected sources exclusively from TSDCD for sources with $F_{low}$ at 5 Hz, 10 Hz, 15 Hz, 20 Hz, and 30 Hz.
FPR is equivalent to the False Alarm Rate (FAR), a common measure in GW detection. FAR quantifies how often a GW detector produces a false detection (a false positive) when there is no actual GW signal present (Morrás et al., 2023). This rate is calculated per unit of data length. Therefore, the term FAR will be used instead.
FAR was computed for each $F_{low}$ for the entire mock data. Figure 9 depicts FAR as a function of $MSNR$, where in the top panel, TSDCD results are compared to SSDD results from WTM1. The general trend is consistent with the previous results in WTM1, where FAR decreases as $MSNR$ increases. As shown in the figure, FAR values have significantly improved, especially for sources with lower $MSNR$, transitioning from 0.712, 0.535, 0.050, 0.014, and 0.001 to 0.526, 0.296, 0.006, 0.004, and 0.0 for $MSNR$ ranges 4-5, 5-6, 6-7, 7-8, and >8, respectively. In the bottom panel, FAR values are depicted for each $F_{low}$. The overall performance is nearly identical across different $F_{low}$ settings, except for the case of $F_{low}$ at 5 Hz. This discrepancy aligns with the accuracy results presented earlier.
In terms of $MSNR$, chirp mass $M$, and luminosity distance $D_{L}$ for undetected sources, Figures 10, 11, and 12 compare the total number of undetected sources between SSDD and TSDCD. Additionally, TSDCD is examined individually for all five different $F_{low}$ values. The right panel in Figure 10 demonstrates a significant improvement across all $MSNR$ ranges, particularly noticeable at relatively higher $MSNR$ 5-8 and greater, and noticeable decrease within $MSNR$ 4-5 where TSDCD exhibits a reduction of 100 undetected sources compared to SSDD. The left panel illustrates variations in the total number of undetected sources across our five $F_{low}$ settings, primarily within the $MSNR$ range of 4-5. Specifically, a total of 548, 531, 549, 562 and 594 undetected sources at $F_{low}$ settings of 5 Hz, 10 Hz, 15 Hz, 20 Hz, and 30 Hz is observed within $MSNR$ range 4-5, respectively.
Similarly, in terms of chirp mass $M$, the right panel in Figure 11 demonstrates a significant decrease in the total number of undetected sources in TSDCD, particularly among sources with chirp masses $M$ in the range of 15-20 $M_{\odot}$. In TSDCD alone, count of undetected sources of different $F_{low}$ varies among all chirp mass $M$ values. For $F_{low}$ at 10 Hz, the minimum count observed is 240, 98, 46,55 and 34 in the chirp mass $M$ ranges of 15-20, 20-25, 25-30,30-35 and 35-40 $M_{\odot}$, respectively. For the chirp mass $M$ ranges 40-45 and 45-50, the minimum count observed are 14 and 7 for $F_{low}$ of 20 Hz and 15 Hz, respectively. Regarding the $D_{L}$, the right panel in Figure 12 shows a substantial reduction in the count of undetected sources in TSDCD, especially for sources with $D_{L}$ in the range of 10-30 Gpc. In the left panel, no significant variation is observed across different $F_{low}$ settings. In summary, both SSDD and TSDCD exhibit a higher detection rate for sources with higher $MSNR$, chirp mass, and shorter $D_{L}$, as indicated by smaller FAR values. Notably, TSDCD demonstrates a significant improvement over SSDD. It’s worth noting that the same source may exhibit different $MSNR$ values on each sub-detector, yet TSDCD consistently outperforms SSDD across these variations.
For a more qualitative check, Table 3 shows the maximum $D_{L}$ ($D_{L_{\text{Max}}}$) in Gpc, minimum $MSNR$ ($MSNR_{\text{Min}}$), and the minimum chirp mass $M_{\text{Min}}$ in $M_{\odot}$, along with the count of detected sources for all $F_{\text{low}}$ out of the total 3,610 sources. The ResNet model successfully detected sources at a distance of 86.601 Gpc, with an average $MSNR$ of 3.9 (averaged $MSNR$ over the three detectors) and a chirp mass of 13.632 at 5 Hz. At 10 Hz, 15 Hz, 20 Hz, and 30 Hz, ResNet successfully identified sources with average $MSNR$ values of 4.031, 4.033, 4.023, and 4.001, along with chirp masses of 20.320, 15.630, 14.201, and 17.532, respectively.
$F_{\text{low}}$ $D_{L_{\text{Max}}}$ $MSNR_{Min}$ $M_{Min}$ Total
5 Hz 86.601 3.900 13.632 3053
10 Hz 90.255 4.031 20.320 3069
15 Hz 67.553 4.033 15.630 3056
20 Hz 67.553 4.023 14.201 3042
30 Hz 67.553 4.001 17.532 3009
Table 3: Maximum $D_{L}$ ($D_{L_{\text{Max}}}$) in Gpc, minimum $MSNR$ ($MSNR_{\text{Min}}$), minimum chirp mass ($M_{\text{Min}}$) in $M_{\odot}$, and count of detected sources for all $F_{\text{low}}$ out of 3,610 total sources.
In terms of computational efficiency, we concurrently processed a total of 25 hours, four minutes, and 16 seconds of data from each detector in parallel. Utilizing the same configuration as in WTM1 (Core i7 MacBook Pro with 16 GB memory, 2667 MHz DDR4, and 2.6 GHz processor), the entire data scan was completed in 22 minutes, averaging 4.7 minutes for each hour. The processing time remains consistent with that achieved on SSDD. The key difference lies in simultaneously reading data from all three sub-detectors, ensuring the computational cost remains unchanged. Hence, utilizing TSDCD is suitable for near-real-time detection and could be further enhanced with a more powerful setup.
6 Conclusion
In this study, we conducted a comparative analysis of the efficiency in detecting BBH sources using data from all three sub-detectors of ET simultaneously versus using data from a single sub-detector. The comparison was specifically focused on cases with $F_{\text{low}}$ set to 30 Hz. Furthermore, we explored TSDCD’s performance across five different $F_{\text{low}}$ settings (5 Hz, 10 Hz, 15 Hz, 20 Hz, and 30 Hz), employing the same $MSNR$ ranges as in our previous study (4-5, 5-6, 6-7, 7-8, and >8). Utilizing the ResNet model, which exhibited superior performance in our previous study, TSDCD demonstrated a significant enhancement in detection accuracy compared to SSDD. The accuracy improved from $60\%$, $60.5\%$, $84.5\%$, $94.5\%$, and $98.5\%$ to $78.5\%$, $84\%$, $99.5\%$, $100\%$, and $100\%$ for sources with $MSNR$ ranges of 4-5, 5-6, 6-7, 7-8, and >8, respectively. The results indicate a substantial accuracy improvement for lower $MSNR$ ranges: 4-5, 5-6, and 6-7, with gains of $18.5\%$, $24.5\%$, and $13\%$ respectively. Additionally, there is a significant improvement of $5.5\%$ and $1.5\%$ for higher $MSNR$ ranges: 7-8 and >8. In qualitative evaluation, the ResNet model detected sources at 86.601 Gpc with an average $MSNR$ of 3.9 (averaged over the three detectors) and a chirp mass of 13.632 at 5 Hz. The complete scan of 25 hours of data (for each sub-detector) took 22 minutes, which equates to 4.7 minutes for each hour. This makes the use of TSDCD appropriate for near-real-time detection and can be greatly improved using a more powerful setup. Future work will focus on including binary neutron stars (BNS) and verifying the performance on real data from LIGO, VIRGO, and KAGRA.
This research was funded in part by National Science Centre, Poland (UMO-2023/49/N/ST9/01287). Additionally, This work has been supported in part by the European Union’s Horizon 2020 research and innovation programme under grant agreement No 952480 (DarkWave project), and from the International Research Agenda Programme AstroCeNT (MAB/2018/7) funded by the Foundation for Polish Science (FNP) from the European Regional Development Fund. The authors thank Prof. Piotr Gawron for his helpful discussions and reviews.